# RewardDS: Privacy-Preserving Fine-Tuning for Large Language Models via Reward Driven Data Synthesis

### v1

@@ Appendix F Extension to More LLM Backbones
As shown in Table 3, RewardDS outperforms other baselines regardless of whether Llama-2-7B-chat-hf or Qwen2.5-14B-Instruct is used as the LLM backbone. This strongly demonstrates that our method is consistently effective, regardless of the LLM backbone. It is worth noting that although Qwen2.5-14B-Instruct has a larger number of parameters compared to Llama-2-7B-chat-hf, our method performs better on the Llama-2-7B-chat-hf model. This is likely due to the use of LoRA fine-tuning on Qwen2.5-14B-Instruct, rather than full-parameter fine-tuning. We believe that applying full-parameter fine-tuning to the Qwen2.5-14B-Instruct model would lead to better performance.
Appendix G Case studies
Here, we provide an example demonstrating how our method refines synthetic data to improve its quality. As shown in Figure 7, the initial synthetic sample contains noise, with redundant and meaningless content highlighted in red. Directly using these synthetic sample will do harm to the fine-tuning process of target LLM. After refinement by RewardDS, the response becomes more coherent and informative, as highlighted in green. Then these refined synthetic sample can be used to fine-tune target LLM for domain-specific tasks.
Figure 7: A synthetic sample refined by our RewardDS on the Medical QA task. Text highlighted with a green background indicates the good and correct response part. Text highlighted with a red background denotes the bad and incorrect response part.
Appendix H Prompt Template Details
H.1 Sampling Queries
Prompt template shown in Figure 8 instructs GPT to act as a data creator by generating a new question similar to given private data from three private datasets. GPT synthetic structured task instructions that align with previous patterns for the subsequently model fine-tuning.
H.2 Sampling Response
Figure 9, 10 and 11 show the prompt templates we employed to sample responses from Medical QA, Financial QA and Code Generation datasets, respectively.
H.3 Generate Feedback
Prompt templates shown in Figure 12, 13 and 14 use LLM-generated feedback to evaluate the strength of chosen responses and the weakness of rejected ones from generation proxy model. The prompt templates are respectively used for Medical QA, Financial QA and Code Generation datasets.
H.4 Refine Synthetic Data

### v2

@@ Appendix H More Analysis of RewardDS Design
Appendix I More Hyperparameter Analysis
In this section, we analyze the other hyperparameters of our method, including the number of folds $k$ and the number of candidate responses $N$, for the medical QA, financial QA and code generation tasks.
As described in Alg.˜1, the number of folds $k$ controls how much of the synthetic data is considered clean. As shown in Figure 9, $k=6$ yields the best performance on the medical QA task. For the financial QA and code generation tasks, the optimal values are $k=5$ (Figure 10) and $k=8$ (Figure 11), respectively. Larger $k$ values lead to stricter filtering, excluding more data, which may cause overfitting on smaller subsets and degrade performance.
As for the number of candidate responses $N$, a larger $N$ increases the likelihood of selecting higher-quality responses but also incurs greater computational cost. As shown in Figure 9, increasing $N$ from 1 to 3 leads to significant performance gains, while further increments yield only marginal improvements. Therefore, we set $N=3$ for the medical QA task. For the financial QA and code generation tasks, we choose $N=2$ (Figure 10) and $N=3$ (Figure 11), respectively.
Figure 9: Performance of RewardDS with different numbers of folds ($k$) and candidate responses (NN) on the dev set for the medical QA task.
Figure 10: Performance of RewardDS with different numbers of folds ($k$) and candidate responses (NN) on the dev set for the financial QA task.
Figure 11: Performance of RewardDS with different numbers of folds ($k$) and candidate responses (NN) on the dev set for the code generation task.
Appendix J Prompt Template Details
J.1 Sampling Queries
Prompt template shown in Figure 12 instructs GPT to act as a data creator by generating a new question similar to given private data from three private datasets. GPT synthesizes structured task instructions that align with previous patterns for the subsequent model fine-tuning.
J.2 Sampling Response
Figure 13, 14 and 15 show the prompt templates we employed to sample responses from Medical QA, Financial QA and Code Generation datasets, respectively.
J.3 Generate Feedback
Prompt templates shown in Figure 16, 17 and 18 use LLM-generated feedback to evaluate the strength of chosen responses and the weakness of rejected ones from generation proxy model. The prompt templates are respectively used for Medical QA, Financial QA and Code Generation datasets.
J.4 Refine Synthetic Data